# BlockingPy: approximate nearest neighbours for blocking of records for entity resolution

### v1

@@ 6 Conclusions
The current project is designed to be extensible so that it can accommodate new ANN algorithms and new input data formats. Any contributions are warmly welcome; see https://github.com/ncn-foreigners/BlockingPy/issues for a feature request and bug tracker. Future developments will focus on using semantic similarity as in the semhash package via custom encoders. Moreover, the framework can be extended to cover privacy preserving record linkage without directly sharing personally identifiable information.
@@ Acknowledgements
This software is a part of Tymoteusz Strojny Bachelor’s thesis at Poznań University of Economics and Business, Poland. The BlockingPy is based on the blocking package in R (see https://ncn-foreigners.github.io/blocking/).
@@
During the preparation of this work the author(s) used DeepL and Claude.ai in order to proofread the paper. After using this the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the published article.
@@
Data are publicly available.

### v2

@@ 6 Conclusions
The current project is designed to be extensible so that it can accommodate new ANN algorithms and new input data formats. Any contributions are warmly welcome; see https://github.com/ncn-foreigners/BlockingPy/issues for a feature request and bug tracker.
Future updates will focus on adding functionality for privacy-preserving record linkage without directly sharing personally identifiable information.
@@ Acknowledgements
We thank the reviewers for their comments on this manuscript. This software is a part of Tymoteusz Strojny’s Bachelor’s thesis at Poznań University of Economics and Business, Poland. BlockingPy is based on the blocking package in R (see https://ncn-foreigners.github.io/blocking/).
@@
During the preparation of this work the author(s) used DeepL and Claude.ai to proofread the paper. After using this the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the published article.
@@
The datasets used in Section 4 are included with the BlockingPy package and are also publicly available from their original sources.